OpenStudy (anonymous):

simplify the expression: 64 ÷ 42 – [2(4 – 6)2]?

OpenStudy (anonymous):

64/24-[2(-2)2] 64/24+8 10.666667

OpenStudy (anonymous):

first comes parentheses and then divide and multiply and add

OpenStudy (trojanpoem):

Take it as parts 2 * (4-6 ) *2 = 2 * -2 * 2 = -8 42 + 8 = 50 64/50 = multiply 2/2 = 128/100 = 1.28

OpenStudy (anonymous):

you do PEMDAS